# Survey of Video Diffusion Models: Foundations, Implementations, and Applications

### v1

@@ 2.5.4 Text encoder
In text-conditional visual generation models, the text encoder plays a crucial role in capturing semantic information from input text prompts, directly affecting the generated content. Early text-to-image models used text encoders trained on paired text-image datasets, either trained from scratch(Nichol et al., 2021; Ramesh et al., 2021) or fine-tuned from pre-trained models such as CLIP(Radford et al., 2021). CLIP, which employs contrastive learning to align text and image embedding spaces, enables the text encoder to effectively represent both visual and linguistic semantics after being trained on large-scale multimodal datasets. Once an input text prompt is tokenized and embedded, it serves as a conditioning mechanism for the diffusion model’s generative backbone. CLIP-based text encoders are widely used in text-to-image diffusion models, as shown in Table 1, including DALLE-2(Ramesh et al., 2022), Stable Diffusion(Rombach et al., 2022), DiT(Bao et al., 2023a; Peebles & Xie, 2023), Fit(Lu et al., 2024c), SiT(Ma et al., 2024a), HunyuanDiT(Li et al., 2024j), and Scaling Diffusion (Esser et al., 2024; Labs, 2024). In these models, the text encoder’s parameters are often frozen to reduce computational and memory overhead during training, ensuring that most training resources are allocated to the diffusion process itself.
@@ 3.1 Datasets
In this part, we review the most popular datasets used for video generation, as summarized in table 2. Specifically, we split the datasets into academic and commercial datasets.
@@
Building upon those traditional benchmark datasets, researchers have recently introduced datasets (Sanabria et al., 2018; Zellers et al., 2021; Lee et al., 2021; Chen et al., 2024c; Ju et al., 2024) with a large number of videos. For example, YT-Temporal-180M (Zellers et al., 2021) contains 180M videos collected from YouTube. InternVid (Wang et al., 2023h) takes over 7 million videos lasting nearly 760K hours, yielding 234M video clips accompanied by detailed descriptions of a total of 4.1B words. Also, researchers noticed that video quality, such as resolution, is also important. Inter4k-1k (Stergiou & Poppe, 2023), as a pioneer, introduces 1k videos with a resolution of 4k, enabling the understanding and generation of high-resolution videos.
On the other side, as manually collecting datasets is time-consuming and expensive, researchers started to generate datasets using off-the-shelf models. VidProM (Wang & Yang, 2024) contains 1.67 million unique text-to-video prompts generated by Pika, VideoCraft2 (Chen et al., 2024a), and Text2Video-Zero (Khachatryan et al., 2023).
On the other side, other researchers focus on providing comprehensive video generation benchmark datasets (Huang et al., 2024a; Wang & Yang, 2024) from different perspectives, e.g., temporal consistency, style consistency, semantic consistency, and video quality. VBench takes 100 prompts to evaluate video generation models from different perspectives. FETV (Liu et al., 2023e) reuses the texts from the MSR-VTT test set and WebVid for open-domain text-video pairs, guaranteeing the diversity of prompts. It also contains manually written prompts describing scenarios that are unusual in the real world, and cannot be found in existing text-video datasets.
@@
UCF-101 (Soomro et al., 2012) V, A Real - AVG. 7 s 320x240 - V (13,320) YouTube CC-BY
@@
VATEX (Wang et al., 2019) V, T, A Real Real AVG. 10 s - - T/V (41 k) Youtube CC-BY-4.0
@@
Panda-70M (Chen et al., 2024c) V, T Real Auto Captioning AVG. 9 s 720p - T/V (70 M) - See license
@@ 3.2 Training engineering
Figure 5: The data preprocessing pipeline of Open-Sora Zheng et al. (2024c).
@@
Data preprocessing. Instead of directly using the original data from the dataset, more and more methods  (Zheng et al., 2024c; Li et al., 2024j; Yang et al., 2024j; Blattmann et al., 2023a) have started to select, filter, and enhance video-text paired data before training the model as shown in fig. 5. OpenSora (Zheng et al., 2024c), as a representative open-sourced method, first splits videos into shorter clips using scene detection (Ravi et al., 2025). Later, clips are evaluated using the aesthetic and optical flow scores, and the clips with low scores are removed from training. After that, the qualified clips are captioned using machine captioners (Hong et al., 2024b; Li et al., 2025; Xu et al., 2024d) and a matching score will be employed to evaluate the alignment between clips and their captions. Finally, samples with adequate scores (high aesthetic quality, large video motion, and strong semantic consistency) are used for training the model. The overall pipeline is presented in fig. 5. Similarly, SVD (Blattmann et al., 2023a) filters out samples with low text-video alignment evaluated by CoCa (Yu et al., 2022a), static videos (evaluated by optical flow), and high text overlays.
@@
Training techniques. Video generation models (Blattmann et al., 2023a; Zheng et al., 2024c) are usually initialized with image models and pretrained with image data. Then, the models are pretrained with video data for learning efficient video representations. After that, models are trained using high-quality data to form the final video generation models. However, due to the difficulty of the video generation problem, different training techniques (Blattmann et al., 2023a; Xing et al., 2023; Hong et al., 2023a; Yang et al., 2024j) are proposed. For example, CogVideoX (Yang et al., 2024j) employs a multi-resolution frame pack strategy to enable the generation of videos with varied lengths. And a progressive training strategy is employed to generate videos of different resolutions. On the other side, to scale up the training of video generation models, multi-node parallelization and VRAM-efficient scheduling are utilized Yang et al. (2024j); Zheng et al. (2024c).
@@
Video generation acceleration techniques. Due to the high computational requirements of training and inferencing video generation methods, researchers have developed several acceleration techniques. For example, Open-Sora (Zheng et al., 2024c) employs flash attention (Dao et al., 2022), ZeRO (Rajbhandari et al., 2020), and gradient checkpoint for kernel optimization, hybrid parallelism, and larger batch size training. Moreover, Open-Sora employs efficient STDiT as the vision encoder while preprocessing text and video data for acceleration. On the other side, parameter-efficient fine-tuning (PEFT) (Geng et al., 2024a; Jia et al., 2022), which is widely applied in training large foundation (language and vision-language) models (Li et al., 2025; Touvron et al., 2023a; b), could be used in video generation for training acceleration (Xing et al., 2024b; Pan et al., 2022). We also include a detailed discussion on long video generation acceleration techniques in section 4.5.
@@ 3.3 Evaluation metrics and benchmarking findings
(Semantic Consistency)] [Overall Alignment] [Temporal Quality / Consistency] [Static Quality] ] ]
@@
Existing evaluation metrics include human evaluation and auto evaluation. Human evaluation (Liu et al., 2023e; Huang et al., 2024a) always includes several aspects such as static quality, temporal quality (consistency), overall alignment, and fine-grained alignment (semantic consistency). And human evaluation is always employed to test the alignment between auto-evaluation and human evaluators, such as VBench (Huang et al., 2024a), which uses 16 human evaluations on 16 different dimensions mainly focusing on temporal and semantic consistency. A human-verified metric is STREAM (Kim et al., 2023b), testified under the same dimensions using auto-evaluation and human evaluation.
@@
On the other side, as auto-evaluation (Huang et al., 2024a; b; Fan et al., 2024) is cheaper and faster, it is widely applied to evaluate video quality and conditioned quality (semantic consistency and temporal consistency). Video quality usually only contains video as the input of metrics while conditioned quality also uses other modal inputs such as image, text, or audio to evaluate the generated video.
@@
Video quality consists of static quality and temporal quality. 1) For static quality, AIGCBench (Fan et al., 2024) employs a pre-trained optical flow estimation model, i.e., RAFT (Teed & Deng, 2020), and calculates the mean square of the average flow score between adjacent frames from the generated video for estimating the motion effects as Flow-Square-Mean. Dover (Wu et al., 2023a) is also employed for evaluating the static quality from aesthetic and technical perspectives. Moreover, the number of video frames is used for accessing the quality. On the other side, FETV (Liu et al., 2023e) employs FID (Heusel et al., 2017) and FVD (Unterthiner et al., 2019) metrics. 2) For temporal quality, AIGCBench assesses it by calculating the average similarity between adjacent frames using CLIP (Radford et al., 2021) noted as GenVideo Clip (Adjacent frames). VBench (Huang et al., 2024a) evaluates it by subject, background, temporal flickering, motion, dynamic degree, aesthetic quality, and imaging quality using multiple pre-trained models including DINO, CLIP, Amt, RAFT, and LAION. Bonneel et al. (2015) proposes a warping error for improving the consistency between frames. Except for those methods, STREAM (Kim et al., 2023b) is proposed to avoid the problem of FVD, which has a stronger emphasis on the spatial aspect than the temporal naturalness of video and demonstrates considerable instability and diverges from human evaluations. It evaluates temporal and semantic quality at the same time.
@@
The metrics for evaluating conditioned quality can be categorized into semantic consistency, temporal consistency, and style consistency. 1) In semantic consistency, for image/video-video metrics, AIGCBench (Fan et al., 2024) employs Mean Squared Error (MSE First) and Structural Similarity Index Measure (SSIM First) for evaluating the first generated frames. To further understand how the semantics are preserved across frames, AIGCBench calculates the cosine similarity between the input image and each frame of generated video using CLIP (Image-GenVideo CLIP). Moreover, GenVideo-RefVideo SSIM is also employed to measure the spatial structural similarity of the generated videos to the reference videos by calculating the SSIM (Structural Similarity Index Measure) between the corresponding frames of the generated and reference videos as the referenced video is at hand. For Text-to-Video metrics, AIGCBench employs GenVideo-Text Clip and GenVideo-RefVideo CLIP (Keyframes), calculating the similarity between the input text / referenced video and keyframes from generated videos using CLIP. Similar to AIGCBench, FETV (Liu et al., 2023e) propose to use stronger models, such as CLIP finetuned on MSR-VTT, BLIP, and UMP, and VQA models for evaluating the semantic consistency. Moreover, FETV also introduces a novel QA-based metric (Otter-VQA). This metric generates yes-no questions on key elements from the text prompt via the Vicuna model and calculates the average number of questions that receive a positive answer by feeding the generated (or ground-truth) videos and questions into the Video LLM as the score. VBench (Huang et al., 2024a) evaluates the semantic consistency by objects, humans, color, relationships, scene, appearance, and temporal style using ViCLIP, CLIP, Tag2Text, UMT, and GRiT. 2) For temporal consistency, researchers usually directly employ the temporal quality metrics. Recently, TC-bench Feng et al. (2024c) proposed to focus on the transition of objects, relations, and background for better temporal consistency. 3) For style consistency, VBench employs CLIP and ViCLIP to assess the similarity between generated videos and the style description.
@@
Except for those metrics, VBench (fig. 6) and Video Generation Arena also set up online evaluation platform, providing an accessible leaderboard for researchers.
@@ 3.4 Industry solutions
As shown Backbone column in Table 1, video generation industry has evolved significantly with the adoption of diffusion models that leverage advanced architectures like DiT, UNet, and U-ViT. DiT (Diffusion Transformer) is widely used in models like W.A.L.T(Gupta et al., 2023), MovieGen, Kling, and Mochi 1(Team, 2024a), excelling in capturing long-range dependencies for sharp, coherent visuals in video generation tasks. UNet, utilized in models like Imagen video, Lavie(Wang et al., 2023g), Seine(Chen et al., 2024d), and VLogger(Zhuang et al., 2024), is known for its ability to handle both local and global features, making it ideal for high-resolution video generation. U-ViT (U-Transformer for video) is another backbone seen in models like VideoCrafter1(Chen et al., 2023a) and VideoCrafter2(Chen et al., 2024a), effective at processing both spatial and temporal dependencies for more accurate video generation. These architectures continue to shape the industry’s approach to advanced video generation.
@@
Text encoders, such as T5-XXL, are essential for aligning video with textual descriptions. Based on the column text encoder in the Table 1, models like Imagen video, W.A.L.T.(Gupta et al., 2023), and MovieGen use T5-XXL for precise video-text alignment, enhancing generative capabilities. Meanwhile, models like Lavie(Wang et al., 2023g) and Seine(Chen et al., 2024d) leverage CLIP-based encoders, which offer flexibility for generating both short clips and full-length videos. T5-XXL excels in large-scale generation, while CLIP provides versatility for diverse visual styles and content.
@@
Training datasets such as WebVid-10M(Bain et al., 2021) and Vimeo25M(Huang et al., 2023b) are commonly used for large-scale models like Seine(Chen et al., 2024d), Lavie(Wang et al., 2023g) and Show-1 from Table 1, ensuring diverse outputs. Models including Lumiereand VideoPoet choose to gather video and image data directly from public websites, rather than relying on pre-existing datasets.
Resolution plays a key role in video output quality. As shown in the resolution coulmn in the Table 1, GEN-3 Alpha and Sora support high resolutions like $4090\times 2160$ and $1920\times 1080$, ideal for cinematic production. In contrast, models like Seine(Chen et al., 2024d) and Phenaki(Villegas et al., 2022) offer lower resolutions like $320\times 512$ or $64\times 64$, balancing quality and computational efficiency.

### v2

@@ 2.5.4 Text encoders
In text-conditional visual generation models, the text encoder plays a crucial role in capturing semantic information from input text prompts, directly affecting the generated content. Early text-to-image models used text encoders trained on paired text-image datasets, either trained from scratch(Nichol et al., 2021; Ramesh et al., 2021) or fine-tuned from pre-trained models such as CLIP(Radford et al., 2021). CLIP, which employs contrastive learning to align text and image embedding spaces, enables the text encoder to effectively represent both visual and linguistic semantics after being trained on large-scale multimodal datasets. Once an input text prompt is tokenized and embedded, it serves as a conditioning mechanism for the diffusion model’s generative backbone. CLIP-based text encoders are widely used in text-to-image diffusion models, as shown in Table 1, including DALLE-2(Ramesh et al., 2022), Stable Diffusion(Rombach et al., 2022), DiT(Bao et al., 2023a; Peebles & Xie, 2023), Fit(Lu et al., 2024c), SiT(Ma et al., 2024a), HunyuanDiT(Li et al., 2024l), and Scaling Diffusion (Esser et al., 2024; Labs, 2024). In these models, the text encoder’s parameters are often frozen to reduce computational and memory overhead during training, ensuring that most training resources are allocated to the diffusion process itself.
@@ 3.1 Datasets
In this part, we review the most popular datasets used for video generation, as summarized in table˜2. Specifically, we split the datasets into academic and commercial datasets.
@@
Building upon those traditional benchmark datasets, researchers have recently introduced datasets (Sanabria et al., 2018; Zellers et al., 2021; Lee et al., 2021; Chen et al., 2024d; Ju et al., 2024) with a large number of videos. For example, YT-Temporal-180M (Zellers et al., 2021) contains 180M videos collected from YouTube. InternVid (Wang et al., 2023h) takes over 7 million videos lasting nearly 760K hours, yielding 234M video clips accompanied by detailed descriptions of a total of 4.1B words. Also, researchers noticed that video quality, such as resolution, is also important. Inter4k-1k (Stergiou & Poppe, 2023), as a pioneer, introduces 1k videos with a resolution of 4k, enabling the understanding and generation of high-resolution videos.
On the other side, as manually collecting datasets is time-consuming and expensive, researchers started to generate datasets using off-the-shelf models. VidProM (Wang & Yang, 2024) contains 1.67 million unique text-to-video prompts generated by Pika, VideoCraft2 (Chen et al., 2024b), and Text2Video-Zero (Khachatryan et al., 2023).
On the other side, other researchers focus on providing comprehensive video generation benchmark datasets (Huang et al., 2024b; Wang & Yang, 2024) from different perspectives, e.g., temporal consistency, style consistency, semantic consistency, and video quality. VBench takes 100 prompts to evaluate video generation models from different perspectives. FETV (Liu et al., 2023e) reuses the texts from the MSR-VTT test set and WebVid for open-domain text-video pairs, guaranteeing the diversity of prompts. It also contains manually written prompts describing scenarios that are unusual in the real world, and cannot be found in existing text-video datasets.
@@
UCF-101 (Soomro et al., 2012) V, A Real - AVG. 7 s 320x240 - V (13,320) YouTube $C$-BY
@@
VATEX (Wang et al., 2019) V, T, A Real Real AVG. 10 s - - T/V (41 k) Youtube $C$-BY-4.0
@@
Panda-70M (Chen et al., 2024d) V, T Real Auto Captioning AVG. 9 s 720p - T/V (70 M) - See license
@@ 3.2 Training engineering
Figure 8: The data preprocessing pipeline of Open-Sora (Zheng et al., 2024c).
Figure 9: The data preprocessing pipeline of WAN (Wan et al., 2025). WAN first splits data into image, video, and text data. Then, different stages of filters are applied to the data for further proprocessing. Fianlly, the data is splited into pretraining data and SFT data for imrpoving the quality of generated video data.
@@
Data preprocessing. Instead of directly using the original data from the dataset, more and more methods  (Zheng et al., 2024c; Li et al., 2024l; Yang et al., 2024j; Blattmann et al., 2023a) have started to select, filter, and enhance video-text paired data before training the model as shown in fig.˜8. OpenSora (Zheng et al., 2024c), as a representative open-sourced method, first splits videos into shorter clips using scene detection (Ravi et al., 2025). Later, clips are evaluated using the aesthetic and optical flow scores, and the clips with low scores are removed from training. After that, the qualified clips are captioned using machine captioners (Hong et al., 2024b; Li et al., 2025a; Xu et al., 2024d) and a matching score will be employed to evaluate the alignment between clips and their captions. Finally, samples with adequate scores (high aesthetic quality, large video motion, and strong semantic consistency) are used for training the model. The overall pipeline is presented in fig.˜8. Similarly, SVD (Blattmann et al., 2023a) filters out samples with low text-video alignment evaluated by CoCa (Yu et al., 2022a), static videos (evaluated by optical flow), and high text overlays.
A recent framework, WAN (Wan et al., 2025), as shown in fig.˜9, employs a novel data pipeline which includes four fundamental dimensions for data collection and selection. For pretraining data, it evaluates data from fundamental dimensions (text detection, blur detection, synthetic data detection, and etc.), visual quality (including data clustering and scoring), motion quality (which splits data with optimal motion, medium-quality motion, camera-driven motion, low-quality motion, and shaky camera footage), and visual text data. For post-training data preprocessing, WAN improves the quality of image and video data separately. Similarly, to meet the needs of different stages of pre-training, teamMeta (2024) curates 3 subsets of pre-training data with progressively stricter visual, motion, and content thresholds, which leads to more stable training process. While MovieGen and WAN apply post-training to improve the motion and aesthetic quality of the generated videos by finetuning the pre-trained model on a small finetuning set of selected videos, Cosmos WFM (NVIDIA et al., 2025) is fine-tuned in post training to support diverse Physical AI applications, such as autonomous driving, with dedicated post-training datasets.
(a) Training pipeline of Cosmos (NVIDIA et al., 2025).
(b) Training pipeline of Movie Gen (teamMeta, 2024).
Figure 10: Training pipelines of Cosmos (NVIDIA et al., 2025) and Movie Gen (teamMeta, 2024). Cosmos first pretrains models and then finetunes on different downstream tasks for better performance while Movie Gen finetunes to achieve better quality of generated videos.
@@
Training techniques. Video generation models (Blattmann et al., 2023a; Zheng et al., 2024c) are usually initialized with image models and pretrained with image data. Then, the models are pretrained with video data for learning efficient video representations. After that, models are trained using high-quality data to form the final video generation models. However, due to the difficulty of the video generation problem, different training techniques (Blattmann et al., 2023a; Xing et al., 2023; Hong et al., 2023a; Yang et al., 2024j) are proposed. For example, CogVideoX (Yang et al., 2024j) employs a multi-resolution frame pack strategy to enable the generation of videos with varied lengths. And a progressive training strategy is employed to generate videos of different resolutions. This progressive training strategy is also employed by other works, such as Hunyuan-Video (Kong et al., 2025). On the other side, to scale up the training of video generation models, multi-node parallelization and VRAM-efficient scheduling are utilized (Yang et al., 2024j; Zheng et al., 2024c). Moreover, as post training has shined the lights on vision-language models (Cheng et al., 2025; NVIDIA et al., 2025; Yamaguchi et al., 2025), more and more studies (Wan et al., 2025; teamMeta, 2024) include post-training to further improve the quality of generated video, as shown in fig.˜10. A practical scenario for post-training is to finetune the pretrained to suit the downstream tasks (NVIDIA et al., 2025), such as autonomous driving and robot manipulation.
@@
Video generation acceleration techniques. Due to the high computational requirements of training and inferencing video generation methods, researchers have developed several acceleration techniques. For example, Open-Sora (Zheng et al., 2024c) employs flash attention (Dao et al., 2022), ZeRO (Rajbhandari et al., 2020), and gradient checkpoint for kernel optimization, hybrid parallelism, and larger batch size training. Moreover, Open-Sora employs efficient STDiT as the vision encoder while preprocessing text and video data for acceleration. On the other side, parameter-efficient fine-tuning (PEFT) (Geng et al., 2024a; Jia et al., 2022), which is widely applied in training large foundation (language and vision-language) models (Li et al., 2025a; Touvron et al., 2023a; b), could be used in video generation for training acceleration (Xing et al., 2024b; Pan et al., 2022). We also include a detailed discussion on long video generation acceleration techniques in section˜4.5.
@@ 3.3 Evaluation metrics and benchmarking findings
Metrics
Video-Condition Consistency
Style Consistency
Temporal (Motion) Consistency
Semantic Consistency
Video Quality
Temporal Quality
Static Quality
Human Evaluation
Fine-grained Alignment
(Semantic Consistency)
Overall Alignment
Temporal Quality / Consistency
Static Quality
@@
Existing evaluation metrics include human evaluation and auto evaluation. Human evaluation (Liu et al., 2023e; Huang et al., 2024b) always includes several aspects such as static quality, temporal quality (consistency), overall alignment, and fine-grained alignment (semantic consistency). And human evaluation is always employed to test the alignment between auto-evaluation and human evaluators, such as VBench (Huang et al., 2024b), which uses 16 human evaluations on 16 different dimensions mainly focusing on temporal and semantic consistency. A human-verified metric is STREAM (Kim et al., 2023c), testified under the same dimensions using auto-evaluation and human evaluation.
@@
On the other side, as auto-evaluation (Huang et al., 2024b; c; Fan et al., 2024) is cheaper and faster, it is widely applied to evaluate video quality and conditioned quality (semantic consistency and temporal consistency). Video quality usually only contains video as the input of metrics while conditioned quality also uses other modal inputs such as image, text, or audio to evaluate the generated video.
@@
Video quality consists of static quality and temporal/motion quality. 1) For static quality, AIGCBench (Fan et al., 2024) employs a pre-trained optical flow estimation model, i.e., RAFT (Teed & Deng, 2020), and calculates the mean square of the average flow score between adjacent frames from the generated video for estimating the motion effects as Flow-Square-Mean. Dover (Wu et al., 2023a) is also employed for evaluating the static quality from aesthetic and technical perspectives. Moreover, the number of video frames is used for accessing the quality. On the other side, FETV (Liu et al., 2023e) employs FID (Heusel et al., 2017) and FVD (Unterthiner et al., 2019) metrics. 2) For temporal/motion quality, AIGCBench assesses it by calculating the average similarity between adjacent frames using CLIP (Radford et al., 2021) noted as GenVideo Clip (Adjacent frames). VBench (Huang et al., 2024b) evaluates it by subject, background, temporal flickering, motion, dynamic degree, aesthetic quality, and imaging quality using multiple pre-trained models including DINO, CLIP, Amt, RAFT, and LAION. Bonneel et al. (2015) proposes a warping error for improving the consistency between frames. Except for those methods, STREAM (Kim et al., 2023c) is proposed to avoid the problem of FVD, which has a stronger emphasis on the spatial aspect than the temporal naturalness of video and demonstrates considerable instability and diverges from human evaluations. It evaluates temporal and semantic quality at the same time.
Besides those metrics, more and more researchers start to improving motion consistency and aesthetics of generated videos. 1) For accessing motion smoothness, WAN (Wan et al., 2025) employs Qwen2-VL (Bai et al., 2023) with crafted prompts. Similarly, Movie Gen (teamMeta, 2024) accesses the motion completeness and motion naturalness for better evaluating the quality of generated videos. 2) While the motion quality of video evaluates how smooth the video is, the realness and aesthetics evaluate the model’s ability to generate photorealistic videos with aesthetically pleasing content, lighting, color, style, etc. To access them, Movie Gen recruits a group of human evaluator to measure which of the videos being compared most closely resembles a real video and which of the generated videos has more interesting and compelling content, lighting, color, and camera effects.
@@
The metrics for evaluating conditioned quality can be categorized into semantic consistency, temporal consistency, and style consistency. 1) In semantic consistency, for image/video-video metrics, AIGCBench (Fan et al., 2024) employs Mean Squared Error (MSE First) and Structural Similarity Index Measure (SSIM First) for evaluating the first generated frames. To further understand how the semantics are preserved across frames, AIGCBench calculates the cosine similarity between the input image and each frame of generated video using CLIP (Image-GenVideo CLIP). Moreover, GenVideo-RefVideo SSIM is also employed to measure the spatial structural similarity of the generated videos to the reference videos by calculating the SSIM (Structural Similarity Index Measure) between the corresponding frames of the generated and reference videos as the referenced video is at hand. For Text-to-Video metrics, AIGCBench employs GenVideo-Text Clip and GenVideo-RefVideo CLIP (Keyframes), calculating the similarity between the input text / referenced video and keyframes from generated videos using CLIP. Similar to AIGCBench, FETV (Liu et al., 2023e) propose to use stronger models, such as CLIP finetuned on MSR-VTT, BLIP, and UMP, and VQA models for evaluating the semantic consistency. Moreover, FETV also introduces a novel QA-based metric (Otter-VQA). This metric generates yes-no questions on key elements from the text prompt via the Vicuna model and calculates the average number of questions that receive a positive answer by feeding the generated (or ground-truth) videos and questions into the Video LLM as the score. VBench (Huang et al., 2024b) evaluates the semantic consistency by objects, humans, color, relationships, scene, appearance, and temporal style using ViCLIP, CLIP, Tag2Text, UMT, and GRiT. 2) For temporal consistency, researchers usually directly employ the temporal quality metrics. Recently, TC-bench (Feng et al., 2024c) proposed to focus on the transition of objects, relations, and background for better temporal consistency. 3) For style consistency, VBench employs CLIP and ViCLIP to assess the similarity between generated videos and the style description.
@@
Except for those metrics, VBench (fig.˜11) and Video Generation Arena also set up online evaluation platform, providing an accessible leaderboard for researchers.
@@ 3.4 Industry solutions
As shown Backbone column in Table 1, video generation industry has evolved significantly with the adoption of diffusion models that leverage advanced architectures like DiT, UNet, and U-ViT. DiT (Diffusion Transformer) is widely used in models like W.A.L.T(Gupta et al., 2023), MovieGen, Kling, and Mochi 1(Team, 2024a), excelling in capturing long-range dependencies for sharp, coherent visuals in video generation tasks. UNet, utilized in models like Imagen video, Lavie(Wang et al., 2023g), Seine(Chen et al., 2024e), and VLogger(Zhuang et al., 2024), is known for its ability to handle both local and global features, making it ideal for high-resolution video generation. U-ViT (U-Transformer for video) is another backbone seen in models like VideoCrafter1(Chen et al., 2023a) and VideoCrafter2(Chen et al., 2024b), effective at processing both spatial and temporal dependencies for more accurate video generation. These architectures continue to shape the industry’s approach to advanced video generation.
@@
Text encoders, such as T5-XXL, are essential for aligning video with textual descriptions. Based on the column text encoder in the Table 1, models like Imagen video, W.A.L.T.(Gupta et al., 2023), and MovieGen use T5-XXL for precise video-text alignment, enhancing generative capabilities. Meanwhile, models like Lavie(Wang et al., 2023g) and Seine(Chen et al., 2024e) leverage CLIP-based encoders, which offer flexibility for generating both short clips and full-length videos. T5-XXL excels in large-scale generation, while CLIP provides versatility for diverse visual styles and content.
@@
Training datasets such as WebVid-10M(Bain et al., 2021) and Vimeo25M(Huang et al., 2023b) are commonly used for large-scale models like Seine(Chen et al., 2024e), Lavie(Wang et al., 2023g) and Show-1 from Table 1, ensuring diverse outputs. Models including Lumiereand VideoPoet choose to gather video and image data directly from public websites, rather than relying on pre-existing datasets.
Resolution plays a key role in video output quality. As shown in the resolution coulmn in the Table 1, GEN-3 Alpha and Sora support high resolutions like $4090\times 2160$ and $1920\times 1080$, ideal for cinematic production. In contrast, models like Seine(Chen et al., 2024e) and Phenaki(Villegas et al., 2022) offer lower resolutions like $320\times 512$ or $64\times 64$, balancing quality and computational efficiency.